Solve by the elimination method:
5x + 6y = 17
5x - y = 38

Answer by user_dude2008(1861)

5x + 6y = 17
5x - y = 38


5x + 6y = 17
30x - 6y = 228
--------------
35x = 245


x=7


5(7)-y=38

35-y=38

y=-3

Answer: x=7, y=-3 -----> (7,-3)